### v1

# Mantis: Interleaved Multi-Image Instruction Tuning

Abstract
The recent years have witnessed a great array of large multimodal models (LMMs) to effectively solve single-image vision language tasks. However, their abilities to solve multi-image visual language tasks is yet to be improved. The existing multi-image LMMs (e.g. OpenFlamingo, Emu, Idefics, etc) mostly gain their multi-image ability through pre-training on hundreds of millions of noisy interleaved image-text data from web, which is neither efficient nor effective. In this paper, we aim at building strong multi-image LMMs via instruction tuning with academic-level resources. Therefore, we meticulously construct Mantis-Instruct containing 721K instances from 14 multi-image datasets. We design Mantis-Instruct to cover different multi-image skills like co-reference, reasoning, comparing, temporal understanding. We combine Mantis-Instruct with several single-image visual-language datasets to train our model Mantis to handle any interleaved image-text inputs. We evaluate the trained Mantis on five multi-image benchmarks and eight single-image benchmarks. Though only requiring academic-level resources (i.e. 36 hours on 16xA100-40G), Mantis-8B can achieve state-of-the-art performance on all the multi-image benchmarks and beats the existing best multi-image LMM Idefics2-8B (pre-trained on 140M interleaved image-text data) by an average of 9 absolute points. We observe that Mantis performs equivalently well on the held-in and held-out evaluation benchmarks, which demonstrates its generalization ability. Remarkably, Mantis only lags behind GPT-4V by an average of 2 absolute points across the five benchmarks. We further evaluate Mantis on single-image benchmarks and demonstrate that Mantis can maintain a strong single-image performance on par with CogVLM and Emu2. Our results are particularly encouraging as it shows that low-cost instruction tuning is indeed much more effective than intensive pre-training in terms of building multi-image LMMs. We will release all of our code and data for reproducibility.
Figure 1: Mantis can accept interleaved text-image to perform different tasks like coherence, comparison, logical reasoning, and temporal understanding, etc. Mantis achieves the state-of-the-art performance on the 5 multi-image benchmarks, and preserves abilities for single-image tasks.
1 Introduction
Large Multimodal Models (LMMs) have advanced significantly in recent years. Both closed-source models like GPT-4V [1], Gemini [2], Reka [3], MM1 [4] and open-source models like LLaVA [5, 6], LLaVA-NeXT [7], BLIP [8], CogVLM [9], Idefics [10] have shown strong visual-language understanding and generation capabilities in single-image tasks like VQA [11], TextVQA [12], GQA [13], etc. Despite the significant interest in improving LMMs’ performance on single-image vision-language tasks, relatively less work attempts to improve LMMs’ capability to solve multi-image vision-language tasks. To the best of our knowledge, only a few open-source models (e.g. Idefics [10], OpenFlamingo [14], Emu2 [15]) are optimized to support multi-image inputs.
We argue that multi-image visual ability is also a crucial ability in real-world applications. Specifically, we categorize the ability into four skills including (1) Co-reference: understanding the references like “second image” in the natural language expression and grounding it on the referred image to generate a response. (2) Reasoning: capturing the information across multiple images and reasoning over these multiple pieces to derive the response. (3) Comparison: capturing the nuances and commonalities between several images. (4) Temporal understanding: observing multiple frames of a video or a scene to understand the temporal information like actions, behaviors, interactions, etc. These are the major four skills we want the LMMs to encompass.
The existing multi-image LMMs like Idefics [10], OpenFlamingo [14], Emu2 [15], MM1 [4] heavily rely on pre-training on a massive amount of interleaved image-text documents like MMC4 [16], Obelics [10] crawled from the web. For example, MMC4 contains 500 million examples, while Obelics contains 140 million examples. Such a pre-training procedure would require a massive amount of computation. Recent works like Mementos [17], BLINK [18], MVBench [19] reveal that they are actually incompetent in handling multi-image inputs despite being extensively pre-trained. We conjecture it’s due to the high volume of noise present in the crawled image-text documents.
In this paper, we attempt to build LMMs with academic-level resources to support interleaved multi-image inputs. Instead of pre-training LMMs with massive amounts of data, we resort to instruction tuning on publicly available academic vision-language datasets. Similar to LLaVA [7], we start from a pre-trained language model like LLaMA-3 [20] or Fuyu [21] and a vision transformer encoder from CLIP [22] or SigLIP [23]. We use the same multimodal projector as LLaVA to map the vision embeddings to the text embeddings space to simply concatenate them with the text embeddings. We also define a text-image interleaving format: “…(image {i}: <BOI>image embeddings</EOI>)…”, where <BOI> and</EOI> are the image delimiters.
We create an instruction-tuning dataset Mantis-Instruct to train our model. Mantis-Instruct has a total of 721K instances, consisting of 14 subsets to cover all the multi-image skills. Among the 14 subsets, 10 subsets are from the existing datasets. For example, NLVR2 [24], IconQA [25], etc are used to cover ‘reasoning’ skill; DreamSim [26], Birds-to-Words [27], etc are used to cover ‘comparing’ skill; NExT-QA [28], STAR [29], etc are used to cover ‘temporal’ skill. We additionally curate four new datasets LLaVA-665k-multi, LRV-multi to cover ‘coref’ skill and Contrast-Caption, Multi-VQA to broaden ‘reasoning’ skill. The detail of all the datasets is included in Table 1. All the subsets are unified in a standardized format, where images and texts are fully interleaved across multi-turn conversations and can be put into any order depending on the task. We train Mantis by combining Mantis-Instruct with another 268K single-image vision-language data to balance the multi-image and single-image abilities.
To evaluate the multi-image skills, we adopt four existing benchmarks NLVR2 [24], QBench [30], BLINK [18], MVBench [19] to cover ‘reasoning’, ‘compare’ and ‘temporal’ skills. We manually create a new challenging dataset Mantis-Eval to cover ‘coref’ and ‘reasoning’ skills. We include 12 competitive baselines including single-image LMMs like LLaVA-Next [7], Qwen-VL [31] and multi-image LMMs like Idefics1/2, OpenFlamingo, and Emu2. For single-image LMMs, we merge multiple images horizontally as a single-image input.
By training on 16 x A100-40G for 36 hours, Mantis achieves the state-of-the-art performance on 5 multi-image tasks, excelling on all the multi-image skills. Mantis is able to outperform the strongest baseline Idefics2 [10] by an average of 9 absolute points. On the held-out evaluation set Mantis-Eval, Mantis beats all the other baselines by at least 9.5 points. These results are encouraging because we show that the low-cost instruction tuning on 721K high-quality data can lead to a much better generalization performance than other LMMs (Emu2, Idefics, etc) intensively pre-trained on 140M-500M noisy web data. Mantis also preserves decent performance of single-image tasks and is on par with CogVLM [9] and Emu2 [15]. We also present insights on the architecture selection and inference approaches based on the results in  section 3. We believe Mantis-Instruct can serve as an essential baseline for future studies on multi-image LMMs due to its efficiency. We will release all the code, data, and models to help the community reproduce our results.
2 Mantis: Interleaved Multi-Image Instruction Tuning
2.1 Method
Model Architecture:
LMM architectures are roughly divided into two classes: decoder-only LMM and encoder-decoder LMM. The most popular ones are Fuyu [21] and LLaVA [32]. We modified these two model structure to enable support for interleaved multi-image input. For Fuyu, we preserve the super-resolution feature containing the maximum image size to be 720P. For LLaVA architecture, we also modify the codes to support multi-image inputs.
Image Context Length:
How many images can a multi-image supported model accept? This is actually limited by both the LLM backbone’s max token length and the number of image tokens the vision encoder generated. For Fuyu, since we limit the image size to 720P, each image consumes at most $10241920\times 1080/30^{2}=$ 1024 image tokens. For the LLaVA architecture, each image consumes a fixed $(336/14)^{2}=576$ image tokens. For Llama3 with maximum window size of 8K, we can feed at most 8 and 14 images respectively. This is fairly enough for all the experiments in this paper.
Text-Image Interleaving Format:
A proper text-image interleaving format can make an LLM easier to acquire multi-image understanding and reasoning ability. We contend that a good text-image interleaving format should: (1) mark boundaries between images clearly, and (2) denote the serial number of images. Following this principle, we designed our interleaving format as follows: "(image {i}: <BOI><image><EOI>)", where <BOI> is the begin of image token and <EOI> is the end of image token. <image> is the placeholder for image patches. This format add clear separators between images, and gives LMM serial information of the image through "image {i}". In practice, we set <BOI> and <EOI> to be <Image> and </Image> respectively.
Multi-Image Skills:
Training data decides the skills and knowledge that we expect the model to learn. As mentioned in  section 1, 4 crucial multi-image skills, co-reference, reasoning, comparison, and temporal understanding are considered. We also deliberately curate our Mantis-Instruct to cover all these skills. For example, we reformat LLaVA-665k to LLaVA-665k-multi for co-reference skills, and VIST for temporal understanding. Details are stated in the following subsection.
2.2 Mantis-Instruct: A large scale multiple image question answering dataset
Subsets Skill # Instance # Avg-I # Max-I # Turns LengthT LengthT+I
Multi-Image Reasoning Datasets
LLaVA-665k-multi Coref 313K 2.0 4 21.4 558 1710
LRV-multi Coref 8K 3.5 9 83.4 2234 4251
NLVR2 Reason 86K 2.0 2 2.0 105 1257
IconQA Reason 64K 2.4 6 2.0 71 1454
Contrast-Caption Reason 36K 3.8 8 7.6 871 3067
ImageCoDe Reason 17K 2.4 6 1.0 126 5886
Multi-VQA Reason 5K 4.0 6 11.0 1102 3417
Coinstruct Compare 151K 2.7 4 7.5 314 1620
Dreamsim Compare 16K 3.0 3 2.0 103 1831
Spot-the-Diff Compare 8K 2.0 2 4.0 121 1273
Birds-to-Words Compare 3K 2.0 2 2.0 101 1253
VIST Temporal 7K 20.3 50 9.7 530 12238
NExT-QA Temporal 4K 8.0 8 17.6 572 5180
STAR Temporal 3K 8.0 8 30.2 961 5569
Total - 721K 4.7 50 14.4 555 3584
Single-Image Reasoning Datasets
DVQA Doc 200K 1.0 1 40.0 304 880
DocVQA Doc 39K 1.0 1 2.0 62 638
ChartQA Chart 28K 1.0 1 2.0 66 642
Total - 268K 1.0 1 14.7 144 720
Table 1: The statistics of Mantis-Instruct. We list the number of examples, average images per example, maximum image per example, average turns, average text token lengths, and average text+image token lengths. $\clubsuit$ means the new subsets we constructed in this paper. Here we use Llava’s tokenizer and assume each image occupies 576 image tokens, which is the number of tokens for each image. Here "Avg LengthT" denotes the average number of tokens for the text part, while "Avg LengthT&I" means the average number of tokens of text and images in total.
We construct Mantis-Instruct from multiple publicly available datasets. Detailed statistics are shown in  Table 1. We use a similar dataset format with LLaVA’s, where each data item contains multiple images and multiple turns of QA pairs are gathered. We report the number of examples, the number of images per item, the average conversation turns, and the average length. We describe the task type of each dataset and the processing techniques used in the followings:
LLaVA-665k-multi: Besides the above-mentioned multi-image reasoning tasks, we also manually reformat the LLaVA-655k dataset from LLaVA-1.5 into a "synthetic" multi-image dataset. We randomly merge 2 to 4 original date items of a single image into a multi-image data item. Then add proper image denotation like "For the second image, …" for each question. Then we shuffle all the conversation pairs to form the final merge items. It’s a "synthetic" multi-image dataset because answering each question in the conversation does not require reasoning across images. However, due to the quality and diversity of the LLaVA-665k data, we contend that it’s helpful for avoiding forgetting via replaying and increasing instruction-following ability.
Co-Instruct [33]: Co-Instruct is a dataset curated to compare the quality of 2-4 images in the format of open-ended answer or multi-choice QA format. It’s similar to the tasks in the Qbench [30].
NLVR2 [24]: NLVR2 is a natural language reasoning dataset grounded in images. The task is to determine whether an statement is true or false given the visual contents of 2 images. NLVR2 covers many linguistic phenomena, such as cardinality, existential quantifiers, universal quantifiers, etc.
Dreamsim [26]: DreamSim proposes a dataset to train a model to judge image similarity. We reuse their datasets, and the task is to judge which candidate image is more similar to a reference image.
ImageCoDe [34]: ImageCoDe is multimodal challenge called image retrieval from contextual descriptions, where a multimodal model is required to retrive/select the correct image from a set of 10 minimally contrastive images. For each image, 9 similar images are retrived from Open Images Dataset V6 [35] using CLIP encodings.
Contrast-Caption: Inspired by the paradigm of contrastive learning, we reformat existing image captioning tasks into a multi-image version, where 2 tasks are defined. Given multiple images, the first task we defined is to judge which image matches the provided caption, and the second task is to generate a caption for a denoted image. We randomly select 2 to 8 images along with their captions to form a single item, then apply the template of the 2 tasks defined. We used the images and captions provided by ShareGPT-4V [36] and LAION GPT-4V captions from LVIS [37].
Spot-the-diff [25]: This is a dataset which requires the model to generate the difference description given 2 images. We use the processed version from MIMIC-IT [38].
LRV-multi [39]: This is a single-image dataset used to mitigate hallucinations of LMMs. We process it similarly to LLaVA-665k-multi to get a "synthetic" multi-image dataset.
Birds-to-Words [27]: The task of this dataset is to generate different descriptions and give 2 images of different birds. We directly take the training split for training. Besides, we prompt GPT-3.5-turbo to generate a possible question given the reference answer, so we can get a diverse question pool.
Multi-VQA: This dataset contains multi-image reasoning QA pairs, where each question is generated by GPT-4 based on the captions of the images. Each question is required to involve at least two images. Images and captions are sampled from ShareGPT-4V [36]. Prompt template is in  Table 5.
Video QA: We also include some video question answering datasets, including NExTQA [28], STAR [29], and visual-story-telling [32]. nextqa and star are the processed version from Flipped-VQA [40]. The visual-story-telling are the processed version from MIMIC-IT [38].
Single-image-VQA: Instead of focusing only on multi-image tasks, we claim that it’s necessary to also include some single-image VQA dataset to avoid catastrophic forgetting on single-image tasks. We included some of the single image data from LLaVA-665k as well as DocVQA [41], DVQA [42], and ChartQA [43] to enhance its ability on diagrams and OCR. In practice, since the size of DVQA is too large, we only use 30k output 200k of it for the training.
Figure 2: Illustrations of Mantis-Instruct datasets.
2.3 Training Details
We have trained multiple versions of our models with different dataset-mixing strategies and different backbones. We mainly build three models. Mantis-8B-CLIP and Mantis-8B-SigLIP are based on Llama-3-8B-Instruct [20], where CLIP [22] and SigLIP [23] are used as the visual encoder respectively. Mantis-8B-Fuyu is based on the Adept AI’s Fuyu-8B [21]. Different from the above architecture, there is not visual encoder for the Fuyu model, where each image path is directly mapped to the token embedding space through a linear layer.
For Mantis-8B-CLIP and Mantis-8B-SigLIP, we first pre-train the multimodal projector on LLaVA pre-train data, where learning rate is set to 1e-3, batch size is set to 256. After pre-training, we fine-tune the 2 models on the Mantis-Instruct. For Mantis-8B-Fuyu, we directly fine-tune Fuyu-8B on Mantis-Instruct without pre-training.
During the fine-tuning, we train each model on the data for 1 epoch, with a batch size being 128. The maximum context length is set to 8192. The learning rate is set to 1e-5, which is half of the learning rate used in LLaVA-1.5’s instruction-tuning. We set the warmup ratio to be 0.03 and use a cosine learning rate scheduler. We speed up our training and inference with Flash-attention2 [44].
3 Experiments
In this section, we aim to evaluate Mantis’s ability on both multi-image and single-image vision-language tasks. For multi-image tasks, depending on whether the LMM can handle multiple images, we either concatenate multiple images (‘merge’) horizontally as a single image or feed multiple images as a ‘sequence’ to the model.
3.1 Baselines
For ‘merge’ evaluation, we include BLIP-2-13B [8], InstructBLIP-13B [45], Qwen-VL-Chat [31], LLaVA-1.5-7B [5], LLaVA-1.6-7B [7], Kosmos2 [46], and Fuyu-8B [21]. For ‘sequence’ evaluation, we include OpenFlamingo [14], Emu2-Chat-34B [15], Idefics1 [10], Idefics2 [10], GPT-4V [1]. Additionally, for single-image tasks, we include InstructBLIP-7B-Vicuna [45], Yi-VL-6B [47].
3.2 Evaluation Benchmarks
For multi-image vision-language tasks, we use 2 held-in benchmarks: NLVR2 [24] and Qbench [30]; and 3 held-out benchmarks: Mantis-Eval, BLINK [18], and MVBench [19]. We report detailed statistics of different datasets in  Table 2.
NLVR2 [24] evaluates the model’s abi.lity to conduct logical reasoning across the contents of images. It asks the model to compare the contents of the two given images and judge whether a given statement is correct or not. All questions are multiple-choice. We use the test-public split for evaluation
Qbench [30] is a benchmark evaluating whether LMMs can properly judge and compare the quality of a benchmark. Q-bench aims to evaluate the low-level visual abilities of LMMs, such as judging the image quality. All questions are multiple-choice. In our experiments, we evaluate the Qbench2-A2-pair dev set, where a low-level visual question is asked based on multiple image contents.
Mantis-Eval consists of 217 multiple-image reasoning examples that cover different topics, including size perceptions, weight comparisons, etc. This dataset is carefully curated by our annotators, where images are acquired via Google Search, and then come up with a proper question which requires understanding the contents in the two images well to answer. Mantis-Eval contains both multiple-choice and short-answer questions. We report results in the test split.
BLINK [18] is a benchmark on core visual perception abilities, where most of the tasks can be solved by humans “within a blink” (e.g., relative depth estimation, visual correspondence, forensics detection, and multi-view reasoning). Some questions in the benchmark involved multiple images, such as image similarity comparison. We report results in the validation set of the benchmark.
MVBench [19] is a comprehensive multimodal video understanding benchmark covering 20 challenging video tasks. The questions cannot be solved with a single frame, which necessitates understanding an image sequence to solve the questions in the benchmark. We report results in test split.
Benchmarks Skill Type # Instances # Avg-I # Max-I LenghtT+I
NLVR2 Reason MCQ 6,967 2.0 2 1,190
Q-Bench Compare MCQ 1,000 2.0 2 1,167
Mantis-Eval Reason & Coref MCQ & Short 217 2.5 5 1,455
BLINK Reason MCQ 1,901 1.9 4 1,129
MVBench Temporal MCQ 4,000 8.0 8 46
Table 2: Statistics of the evaluation dataset. NLVR2 and Mantis-Eval are held-in dataset, while the rest are held-out. For MVBench, we extract 8 frames per video for the evaluation.
Models Size Format NLVR2 Q-Bench Mantis-Eval BLINK MVBench Avg
GPT-4V - sequence 88.80 76.52 62.67 51.14 43.50 64.5
Open Source Models
Random - - 48.93 40.20 23.04 38.09 27.30 35.5
Kosmos2 1.6B merge 49.00 35.10 30.41 37.50 21.62 34.7
LLaVA-1.5 7B merge 53.88 49.32 31.34 37.13 36.00 41.5
LLaVA-1.6 7B merge 58.88 54.80 45.62 39.55 40.90 47.9
[…]
Fuyu 8B merge 51.10 49.15 27.19 36.59 30.20 38.8
[…]
InstructBLIP 13B merge 60.26 44.30 45.62 42.24 32.50 44.9
CogVLM 17B merge 58.58 53.20 45.16 41.54 37.30 47.1
[…]
OpenFlamingo 9B sequence 36.41 19.60 12.44 39.18 7.90 23.1
Idefics1 9B sequence 54.63 30.60 28.11 24.69 26.42 32.9
Emu2-Chat 37B sequence 58.16 50.05 37.79 36.20 39.72 44.4
Mantis-Fuyu 8B sequence 72.41 54.10 37.79 39.87 39.12 48.6
Mantis-CLIP 8B sequence 84.66 66.00 55.76 47.06 48.30 60.3
Mantis-SigLIP 8B sequence 87.43 69.90 59.45 46.35 50.15 62.6
$\Delta$ over SoTA - - +0.56 +12.90 +9.68 +1.88 +8.00 +9.1
Table 3: Evaluation results on multi-image benchmarks. The highest score of open-source models for each tasks is bolded. The highest score of the open-source baseline model is underscored. $\Delta$ is the performance gains of Mantis compared to the best baseline model for each task.
3.3 Results on Multi-Image Tasks
Model Size TextVQA VQA MMB MMMU OKVQA SQA MathVista Avg
InstructBLIP 7B 33.6 75.2 38.3 30.6 45.2 70.6 24.4 48.9
[…]
LLaVA-1.5 7B 58.2 76.6 64.8 35.3 53.4 70.4 25.6 59.8
Qwen-VL-Chat 7B 63.8 78.2 61.8 35.9 56.6 68.2 15.5 60.8
[…]
Idefics1 9B 39.3 68.8 45.3 32.5 50.4 51.6 21.1 48.0
OpenFlamingo 9B 46.3 58.0 32.4 28.7 51.4 45.7 18.6 43.8
[…]
Table 4: Evaluation results of Mantis on various single-image tasks. The full name of the abbreviations in the table are listed here. TextVQA: TextVQA validation set; VQA: VQA-v2 validation set; MMB: MMBench-English test set; SQA: ScienceQA-Image validation set. The highest score for each task is bolded, while the second highest is underscored.
We report our results on the multi-image task in  Table 3. We put the GPT-4V’s results at the top of the table as it’s a closed-source LMM. We put comparison results between Mantis with other powerful open-source LMMs, along with the insights derived in the following paragraphs.
Mantis performs well on held-in evaluation Mantis-8B-SigLIP achieves $87.43$ and $69.90$ on NLVR2 and Q-Bench respectively, marking it the best open-source LMM on these 2 tasks. When compared to Idefics2, which also trains on NLVR2 training set, our result is marginally better. It’s worth noticing that the Mantis-8B-SigLIP is only 1.37 behind GPT-4V on NLVR2. The highest scores on these 2 benchmarks demonstrate that Mantis has gained strong ability in multi-image logical reasoning and low-level vision understanding.
Mantis generalizes well on held-out evaluation Generalization of held-out multi-image tasks is a more valuable metric. Our evaluation results on the 3 held-out benchmarks, Mantis-Eval, BLINK (val), and MVBench also demonstrate our strong generalization ability. Mantis-8B-SigLIP achieves $59.45$, $46.35$, and $50.15$ respectively on the 3 held-out tasks, surpassing all other open-source baselines. The gain over existing strongest baseline are particulouly tremendous on Mantis-Eval and MVBench.
Decoder-only LMM vs encoder-decoder LMM Unlike many other LMMs like Flamingo, LLaVA, and BLIP, Fuyu is a decoder-only LMM without a vision encoder. A big advantage of this practice is that it supports super-resolution, so the image will not be resized to fit the vision encoder. So which architecture is better? To investigate deeper, we also train an additional Mantis-8B-Fuyu as a comparison. Results in  Table 3 show that the performance of Mantis-8B-Fuyu is worse than Mantis-8B-CLIP and Mantis-8B-SigLIP where a vision encoder exists, even after fine-tuning on the same set of data. We thus conclude vision encoder should be necessary. This is also consistent with our intuition, where people see more, but think/write/talk less. In short, Encoder-Decoder LMM might be a better choice for LMM architecture.
Vision encoder matters As stated above, the Encoder-Decoder is a superior architecture compared to the decoder-only LMM. Following this, the quality of the vision encoder can also have big effects on the final LMM performance. We include 2 versions of Mantis, which is trained with CLIP [22] and SigLIP [23] respectively. And Mantis-8B-SIGIP can achieve generally better performance than the Mantis-8B-CLIP. This shows that SigLIP is a better vision encoder for LMM development.
‘Merge’ vs ‘Sequence’ input The ‘merge’ evaluation is an important baseline to evaluate the multi-image understanding ability of those LMMs that can accept one image at a time. If the results are good by merging images, then interleaved text/image instruction tuning will seem unnecessary. However, it’s not the truth. LLaVA-v1.6’s dynamic image resolution feature can divide a single image into multiple tiles, processing like multiple images [7]. But it still behaves badly on multi-image tasks, only getting $45.62$ and $39.55$ on the Mantis-Eval and BLINK benchmark. Though the results are better than LLaVA-v1.5, which does not have the dynamic image resolution feature, it’s still limited by the image resolution and cannot recognize separate images well.
3.4 Results on Single-Image Tasks
We also evaluate Mantis-8B-CLIP and Mantis-8B-SigLIP on various single-image tasks, including TextVQA [48], VQA-v2 [49], MMBench [50], MMMU [51], etc. Results are shown in  Table 4. All the evaluations are conducted with the help of LMMs-Eval [52] tool.
Compared to a set of previous famous LMMs, Mantis-8B-SigLIP gets significant improvements on MMBench-English, MMMU, and ScienceQA benchmarks. Solving problems in these tasks requires not only the visual perception ability but also the good reasoning ability of the LLM backbone. Considering that we are using the powerful LLaMA-3 as the backbone LLM, these improvements are thus expected. However, it’s also worth noticing the performance degeneration on some tasks after introducing the multi-image ability. Compared to LLaVA-v1.5-7B, Mantis-8B-CLIP’s performance on TextVQA, VQA-v2 drops $1.8$ and $3.6$ respectively. However, this decrease is acceptable as a trade-off for enhancing multi-image ability.
3.5 Case Studies
Figure 3: Case Study of Mantis. The Mantis model is able to understand multiple images and complete the given task
As shown in  Figure 3, we present 2 cases where Mantis is doing do compared to LLaVA-1.5, which does not have multi-image ability, along with Emu-2 and GPT-4V, which also have multi-image ability. Mantis can understand multiple images and generate a more accurate answer.
4 Related Works
4.1 Large Multimodal Models
Large Multimodal Models (LMMs) are mostly denoted as large language models that can understand multiple modalities beside the human language. Some works aim to fuse image, audio, video, or more modalities with language [53, 54], while more works mainly focus on better combining vision knowledge with the language [55, 14, 45]. BLIP-2 has curated a large-scale image captioning dataset, bootstrapping a language model along with a vision encoder to be a powerful multimodal model [8]. Following this, LLaVA further proposes a cheaper approach to train a powerful LMM through visual instruction tuning [32]. LLaVA-Next further optimizes single-image performance, with a cost of an increasing number of image tokens per token [7], consuming more than 2k for each image, which is about 4 times of the original LLaVA model. Later LMMs like QwenVL [31], CogVLM [9], Yi-VL [47], etc. all follow a similar architecture of LLaVA. However, real-world visual problems are way more complex and usually require a sequence of images to describe the situation. Although better performance is achieved in the single-image scenario, it’s doubtful whether it’s worth the cost.
4.2 Multi-Image Supported Large Multimodal Models
Previous works have also noticed the importance of multi-image ability. Deepmind’s close-sourced Flamingo [55] focuses on optimizing in-context learning for a sequence of image/text pairs but lacking free-from image-text interleaved training examples. Kosmos2 [46], Fuyu [21] both support text-image interleaved format given the model structure, but they did not optimize in multi-image reasoning, and the released inference codes also do not support multi-image as input. Emu2 [15] is a generative multimodal model that supports interleaved text-image inputs, as well as generating both images and texts. However, it’s mainly focused on in-context examples like Flamingo and auto-regressive image generation, instead of the 4 skills mentioned in  section 1 Another common variant of LMMs that can accept multiple images as input is video understanding models, such as Video-LLaMA [56]. However, as shown in the MVBench benchmark, it is also shown that Video-LLaMA behaves worse than LMMs that can only accept one image as input [19], which raises doubts about its multi-image understanding ability. Different from these works, Mantis aims to optimize the multi-image ability of LMMs guided by the 4 defined skills, co-reference, reasoning, comparison, and temporal understanding, thus equipping the model with better multi-image ability.
4.3 Multi-Image Evaluation
While many benchmarks have been developed for single-image tasks, such as TextVQA [48], VQA-V2 [49], MMBench [50], and so on, the multi-image evaluation did not capture much of the people’s attention until recently. Difference description is the first task which large-scale data and good evaluation benchmarks are developed for, including works like Spot-the-Diff [25], Birds-to-Words [27], etc. They are usually evaluated by n-gram metrics like BLEU, and ROUGE, and only involve two images. The later NLVR2 [24] further includes logical reasoning across 2 images into evaluation, focusing on a higher-level of cognition ability. Q-bench [30] and BLINK [18] are both proposed for low-level multi-image vision tasks, where human easily do but model struggles. Another set of multi-image evaluation comes from video or image sequence evaluation, including Mementos [17], MVBench [19], etc. In this case, model usually need to take more than 2 images for answering.
5 Conclusion
We propose Mantis, a new large multimodal model family designed to be able to accept interleaved text-image inputs. Mantis is fine-tuned on our Mantis-Instruct, a text-image interleaved dataset consisting of 721K examples, covering 4 crucial skills, including co-reference, reasoning, comparison, and temporal understanding, for multi-image tasks. We also propose Mantis-Eval, a high-quality multi-image evaluation benchmark consisting of 217 examples with an average of 2.5 images per example. Comprehensive experiments show that Mantis models effectively acquire the 4 multi-image skills and achieves state-of-the-art performance on 5 multi-image evaluation benchmarks, and preserve decent single-image reasoning performance. We shared insights about the model architecture and inference approaches based on the experiment results. Extending the image sequence context length and compressing the number of tokens per tokens efficiently are regarded as our future work.
Appendix A Appendix / supplemental material
A.1 Heuristics for data curation
Similar to visual instruction tuning [5], data curation is an important step to ensure good performance. During the curation, we have found some heuristics that can improve the performance. We list them below.
  1. 1.
    Conflicts between short answers and long answers. Many datasets we collected are domain-specific, and some of them are classification tasks like NLVR2 [24] and DreamSim [26]. For these datasets, we convert them into multiple-choice QA format to avoid conflicts between repeated short-answers and some long answers from other datasets.
  2. 2.
    Addition of image denotation in the text for each question. We manually add image denotations like "the first image", "image 2", "the right image", etc. to each question in our dataset to make sure the question involved with multiple images is clear to answer.
  3. 3.
    Positions of the "<image>" placeholder. For each item with multiple images, we write some rules to randomly put the image placeholders at the beginning of the first question or the end of the first question. This technique is also used by the curation LLaVA-665k [6].
Appendix B Prompt Template
Please generate 10 independent QA pairs. Each question shall involve at least 2 images to answer.
Try to cover different ability like reasoning, planning, common sense understanding, etc. Be
creative with your questions and **make sure the answers require integration of information
from multiple images**. Use "image i" to refer to the i-th image in your questions.
Output format:
Question: First question?
Answer: The answer to the first question.
Question: Second question?
Answer: The answer to the second question.
Table 5: Propmt template used to curate the Synthetic-GPT4 subset.